\providecommand{\href}[2]{#2}\begingroup\raggedright\begin{thebibliography}{100}

\bibitem{Duff:1987qa}
M.~J. Duff, ``{Supermembranes: The First Fifteen Weeks},''
  \href{http://dx.doi.org/10.1088/0264-9381/5/1/023}{{\em Class. Quant. Grav.}
  {\bf 5} (1988)  189}.

\bibitem{Bergshoeff:1987dh}
E.~Bergshoeff, M.~J. Duff, C.~N. Pope, and E.~Sezgin, ``{Supersymmetric
  Supermembrane Vacua and Singletons},''
  \href{http://dx.doi.org/10.1016/0370-2693(87)91465-1}{{\em Phys. Lett. B}
  {\bf 199} (1987)  69--74}.

\bibitem{Blencowe:1987bn}
M.~P. Blencowe and M.~J. Duff, ``{Supersingletons},''
  \href{http://dx.doi.org/10.1016/0370-2693(88)90544-8}{{\em Phys. Lett. B}
  {\bf 203} (1988)  229--236}.

\bibitem{Duff:1988st}
M.~J. Duff and C.~Sutton, ``{The Membrane at the End of the Universe},'' {\em
  New Sci.} {\bf 118} (1988)  67--71.

\bibitem{Nicolai:1988ek}
H.~Nicolai, E.~Sezgin, and Y.~Tanii, ``{Conformally Invariant Supersymmetric
  Field Theories on $S^p \times S^1$ and Super p-branes},''
  \href{http://dx.doi.org/10.1016/0550-3213(88)90077-6}{{\em Nucl. Phys. B}
  {\bf 305} (1988)  483--496}.

\bibitem{Bergshoeff:1988uc}
E.~Bergshoeff, M.~J. Duff, C.~N. Pope, and E.~Sezgin, ``{Compactifications of
  the Eleven-Dimensional Supermembrane},''
  \href{http://dx.doi.org/10.1016/0370-2693(89)91053-8}{{\em Phys. Lett. B}
  {\bf 224} (1989)  71--78}.

\bibitem{Duff:1989ez}
M.~J. Duff, C.~N. Pope, and E.~Sezgin, ``{A Stable Supermembrane Vacuum With a
  Discrete Spectrum},''
  \href{http://dx.doi.org/10.1016/0370-2693(89)90575-3}{{\em Phys. Lett. B}
  {\bf 225} (1989)  319--324}.

\bibitem{Bergshoeff:1989av}
E.~Bergshoeff, E.~Sezgin, and Y.~Tanii, ``{Stress Tensor Commutators and
  Schwinger Terms in Singleton Theories},''
  \href{http://dx.doi.org/10.1142/S0217751X90001550}{{\em Int. J. Mod. Phys. A}
  {\bf 5} (1990)  3599--3616}.

\bibitem{Duff:1989aj}
M.~J. Duff, ``{Classical and Quantum Supermembranes},''
  \href{http://dx.doi.org/10.1088/0264-9381/6/11/012}{{\em Class. Quant. Grav.}
  {\bf 6} (1989)  1577--1598}.

\bibitem{Claus:1998mw}
P.~Claus, R.~Kallosh, J.~Kumar, P.~K. Townsend, and A.~Van~Proeyen,
  ``{Conformal theory of M2, D3, M5 and D1-branes + D5-branes},''
  \href{http://dx.doi.org/10.1088/1126-6708/1998/06/004}{{\em JHEP} {\bf 06}
  (1998)  004}, \href{http://arxiv.org/abs/hep-th/9801206}{{\tt
  arXiv:hep-th/9801206}}.

\bibitem{Bergshoeff:1987cm}
E.~Bergshoeff, E.~Sezgin, and P.~K. Townsend, ``{Supermembranes and
  Eleven-Dimensional Supergravity},''
  \href{http://dx.doi.org/10.1016/0370-2693(87)91272-X}{{\em Phys. Lett. B}
  {\bf 189} (1987)  75--78}.

\bibitem{Dirac:1963ta}
P.~A.~M. Dirac, ``{A Remarkable representation of the 3 + 2 de Sitter group},''
  \href{http://dx.doi.org/10.1063/1.1704016}{{\em J. Math. Phys.} {\bf 4}
  (1963)  901--909}.

\bibitem{Fronsdal:1981gq}
C.~Fronsdal, ``{The Dirac Supermultiplet},''
  \href{http://dx.doi.org/10.1103/PhysRevD.26.1988}{{\em Phys. Rev. D} {\bf 26}
  (1982)  1988}.

\bibitem{Nicolai:1984gb}
H.~Nicolai and E.~Sezgin, ``{Singleton Representations of Osp($N$,4)},''
  \href{http://dx.doi.org/10.1016/0370-2693(84)91488-6}{{\em Phys. Lett. B}
  {\bf 143} (1984)  389--395}.

\bibitem{Nahm:1977tg}
W.~Nahm, ``{Supersymmetries and their Representations},''
  \href{http://dx.doi.org/10.1016/0550-3213(78)90218-3}{{\em Nucl. Phys. B}
  {\bf 135} (1978)  149}.

\bibitem{VanProeyen:1999ni}
A.~Van~Proeyen, ``{Tools for supersymmetry},'' {\em Ann. U. Craiova Phys.} {\bf
  9} (1999) no.~I, 1--48, \href{http://arxiv.org/abs/hep-th/9910030}{{\tt
  arXiv:hep-th/9910030}}.

\bibitem{Gunaydin:1998km}
M.~Gunaydin and D.~Minic, ``{Singletons, doubletons and M theory},''
  \href{http://dx.doi.org/10.1016/S0550-3213(98)00225-9}{{\em Nucl. Phys. B}
  {\bf 523} (1998)  145--157}, \href{http://arxiv.org/abs/hep-th/9802047}{{\tt
  arXiv:hep-th/9802047}}.

\bibitem{Seiberg:1999xz}
N.~Seiberg and E.~Witten, ``{The D1 / D5 system and singular CFT},''
  \href{http://dx.doi.org/10.1088/1126-6708/1999/04/017}{{\em JHEP} {\bf 04}
  (1999)  017}, \href{http://arxiv.org/abs/hep-th/9903224}{{\tt
  arXiv:hep-th/9903224}}.

\bibitem{Batrachenko:2002pu}
A.~Batrachenko, M.~J. Duff, and J.~X. Lu, ``{The Membrane at the end of the (de
  Sitter) universe},''
  \href{http://dx.doi.org/10.1016/j.nuclphysb.2006.09.017}{{\em Nucl. Phys. B}
  {\bf 762} (2007)  95--111}, \href{http://arxiv.org/abs/hep-th/0212186}{{\tt
  arXiv:hep-th/0212186}}.

\bibitem{Horowitz:1991cd}
G.~T. Horowitz and A.~Strominger, ``{Black strings and P-branes},''
  \href{http://dx.doi.org/10.1016/0550-3213(91)90440-9}{{\em Nucl. Phys. B}
  {\bf 360} (1991)  197--209}.

\bibitem{Duff:1991pea}
M.~J. Duff and J.~X. Lu, ``{The Selfdual type IIB superthreebrane},''
  \href{http://dx.doi.org/10.1016/0370-2693(91)90290-7}{{\em Phys. Lett. B}
  {\bf 273} (1991)  409--414}.

\bibitem{Duff:1992hu}
M.~J. Duff and J.~X. Lu, ``{Type II p-branes: The Brane scan revisited},''
  \href{http://dx.doi.org/10.1016/0550-3213(93)90457-Z}{{\em Nucl. Phys. B}
  {\bf 390} (1993)  276--290}, \href{http://arxiv.org/abs/hep-th/9207060}{{\tt
  arXiv:hep-th/9207060}}.

\bibitem{Duff:1994an}
M.~J. Duff, R.~R. Khuri, and J.~X. Lu, ``{String solitons},''
  \href{http://dx.doi.org/10.1016/0370-1573(95)00002-X}{{\em Phys. Rept.} {\bf
  259} (1995)  213--326}, \href{http://arxiv.org/abs/hep-th/9412184}{{\tt
  arXiv:hep-th/9412184}}.

\bibitem{Polchinski:1995mt}
J.~Polchinski, ``{Dirichlet Branes and Ramond-Ramond charges},''
  \href{http://dx.doi.org/10.1103/PhysRevLett.75.4724}{{\em Phys. Rev. Lett.}
  {\bf 75} (1995)  4724--4727}, \href{http://arxiv.org/abs/hep-th/9510017}{{\tt
  arXiv:hep-th/9510017}}.

\bibitem{Polchinski:1996na}
J.~Polchinski, ``{Tasi lectures on D-branes},''
  \href{http://arxiv.org/abs/hep-th/9611050}{{\tt arXiv:hep-th/9611050}}.

\bibitem{Gueven:1992hh}
R.~Gueven, ``{Black p-brane solutions of D = 11 supergravity theory},''
  \href{http://dx.doi.org/10.1016/0370-2693(92)90540-K}{{\em Phys. Lett. B}
  {\bf 276} (1992)  49--55}.

\bibitem{Beck:2017wpm}
S.~Beck, U.~Gran, J.~Gutowski, and G.~Papadopoulos, ``{All Killing
  Superalgebras for Warped AdS Backgrounds},''
  \href{http://dx.doi.org/10.1007/JHEP12(2018)047}{{\em JHEP} {\bf 12} (2018)
  047}, \href{http://arxiv.org/abs/1710.03713}{{\tt arXiv:1710.03713
  [hep-th]}}.

\bibitem{Haupt:2018gap}
A.~S. Haupt, S.~Lautz, and G.~Papadopoulos, ``{A non-existence theorem for $N >
  16$ supersymmetric AdS$_3$ backgrounds},''
  \href{http://dx.doi.org/10.1007/JHEP07(2018)178}{{\em JHEP} {\bf 07} (2018)
  178}, \href{http://arxiv.org/abs/1803.08428}{{\tt arXiv:1803.08428
  [hep-th]}}.

\bibitem{Achucarro:1987nc}
A.~Achucarro, J.~M. Evans, P.~K. Townsend, and D.~L. Wiltshire, ``{Super
  p-Branes},'' \href{http://dx.doi.org/10.1016/0370-2693(87)90896-3}{{\em Phys.
  Lett. B} {\bf 198} (1987)  441--446}.

\bibitem{Gunaydin:1986cs}
M.~Gunaydin, B.~E.~W. Nilsson, G.~Sierra, and P.~K. Townsend, ``{Singletons and
  Superstrings},'' \href{http://dx.doi.org/10.1016/0370-2693(86)90922-6}{{\em
  Phys. Lett. B} {\bf 176} (1986)  45--49}.

\bibitem{Ferrara:1998zt}
S.~Ferrara, M.~Porrati, and A.~Zaffaroni, ``{N=6 supergravity on AdS(5) and the
  SU(2,2/3) superconformal correspondence},''
  \href{http://dx.doi.org/10.1023/A:1007592711262}{{\em Lett. Math. Phys.} {\bf
  47} (1999)  255--263}, \href{http://arxiv.org/abs/hep-th/9810063}{{\tt
  arXiv:hep-th/9810063}}.

\bibitem{Aharony:2015oyb}
O.~Aharony and M.~Evtikhiev, ``{On four dimensional N = 3 superconformal
  theories},'' \href{http://dx.doi.org/10.1007/JHEP04(2016)040}{{\em JHEP} {\bf
  04} (2016)  040}, \href{http://arxiv.org/abs/1512.03524}{{\tt
  arXiv:1512.03524 [hep-th]}}.

\bibitem{Garcia-Etxebarria:2015wns}
l.~Garcia-Etxebarria and D.~Regalado, ``{$ \mathcal{N}=3 $ four dimensional
  field theories},'' \href{http://dx.doi.org/10.1007/JHEP03(2016)083}{{\em
  JHEP} {\bf 03} (2016)  083}, \href{http://arxiv.org/abs/1512.06434}{{\tt
  arXiv:1512.06434 [hep-th]}}.

\bibitem{Aharony:2016kai}
O.~Aharony and Y.~Tachikawa, ``{S-folds and 4d N=3 superconformal field
  theories},'' \href{http://dx.doi.org/10.1007/JHEP06(2016)044}{{\em JHEP} {\bf
  06} (2016)  044}, \href{http://arxiv.org/abs/1602.08638}{{\tt
  arXiv:1602.08638 [hep-th]}}.

\bibitem{Borsten:2018jjm}
L.~Borsten, M.~J. Duff, and A.~Marrani, ``{Twin conformal field theories},''
  \href{http://dx.doi.org/10.1007/JHEP03(2019)112}{{\em JHEP} {\bf 03} (2019)
  112}, \href{http://arxiv.org/abs/1812.11130}{{\tt arXiv:1812.11130
  [hep-th]}}.

\bibitem{Townsend:1987yy}
P.~K. Townsend, ``{Supersymmetric extended solitons},''
  \href{http://dx.doi.org/10.1016/0370-2693(88)90852-0}{{\em Phys. Lett. B}
  {\bf 202} (1988)  53--57}.

\bibitem{Strominger:1990et}
A.~Strominger, ``{Heterotic solitons},''
  \href{http://dx.doi.org/10.1016/0550-3213(90)90599-9}{{\em Nucl. Phys. B}
  {\bf 343} (1990)  167--184}. [Erratum: Nucl.Phys.B 353, 565--565 (1991)].

\bibitem{Duff:1991sz}
M.~J. Duff and J.~X. Lu, ``{A Duality between strings and five-branes},''
  \href{http://dx.doi.org/10.1088/0264-9381/9/1/004}{{\em Class. Quant. Grav.}
  {\bf 9} (1992)  1--16}.

\bibitem{Duff:1991pe}
M.~J. Duff, R.~R. Khuri, and J.~X. Lu, ``{String and five-brane solitons:
  Singular or nonsingular?},''
  \href{http://dx.doi.org/10.1016/0550-3213(92)90025-7}{{\em Nucl. Phys. B}
  {\bf 377} (1992)  281--294}, \href{http://arxiv.org/abs/hep-th/9112023}{{\tt
  arXiv:hep-th/9112023}}.

\bibitem{Callan:1991at}
C.~G. Callan, Jr., J.~A. Harvey, and A.~Strominger, ``{Supersymmetric string
  solitons},'' \href{http://arxiv.org/abs/hep-th/9112030}{{\tt
  arXiv:hep-th/9112030}}.

\bibitem{Gibbons:1993xt}
G.~W. Gibbons, D.~Kastor, L.~A.~J. London, P.~K. Townsend, and J.~H. Traschen,
  ``{Supersymmetric selfgravitating solitons},''
  \href{http://dx.doi.org/10.1016/0550-3213(94)90558-4}{{\em Nucl. Phys. B}
  {\bf 416} (1994)  850--880}, \href{http://arxiv.org/abs/hep-th/9310118}{{\tt
  arXiv:hep-th/9310118}}.

\bibitem{Duff:1987bx}
M.~J. Duff, P.~S. Howe, T.~Inami, and K.~S. Stelle, ``{Superstrings in D=10
  from Supermembranes in D=11},''
  \href{http://dx.doi.org/10.1016/0370-2693(87)91323-2}{{\em Phys. Lett. B}
  {\bf 191} (1987)  70}.

\bibitem{Duff:1990xz}
M.~J. Duff and K.~S. Stelle, ``{Multimembrane solutions of D = 11
  supergravity},'' \href{http://dx.doi.org/10.1016/0370-2693(91)91371-2}{{\em
  Phys. Lett. B} {\bf 253} (1991)  113--118}.

\bibitem{Duff:1994fg}
M.~J. Duff, G.~W. Gibbons, and P.~K. Townsend, ``{Macroscopic superstrings as
  interpolating solitons},''
  \href{http://dx.doi.org/10.1016/0370-2693(94)91260-2}{{\em Phys. Lett. B}
  {\bf 332} (1994)  321--328}, \href{http://arxiv.org/abs/hep-th/9405124}{{\tt
  arXiv:hep-th/9405124}}.

\bibitem{Hull:1994ys}
C.~M. Hull and P.~K. Townsend, ``{Unity of superstring dualities},''
  \href{http://dx.doi.org/10.1016/0550-3213(94)00559-W}{{\em Nucl. Phys. B}
  {\bf 438} (1995)  109--137}, \href{http://arxiv.org/abs/hep-th/9410167}{{\tt
  arXiv:hep-th/9410167}}.

\bibitem{Witten:1995ex}
E.~Witten, ``{String theory dynamics in various dimensions},''
  \href{http://dx.doi.org/10.1016/0550-3213(95)00158-O}{{\em Nucl. Phys. B}
  {\bf 443} (1995)  85--126}, \href{http://arxiv.org/abs/hep-th/9503124}{{\tt
  arXiv:hep-th/9503124}}.

\bibitem{Duff:1996aw}
M.~J. Duff, ``{M theory (The Theory formerly known as strings)},''
  \href{http://dx.doi.org/10.1142/S0217751X96002583}{{\em Int. J. Mod. Phys. A}
  {\bf 11} (1996)  5623--5642}, \href{http://arxiv.org/abs/hep-th/9608117}{{\tt
  arXiv:hep-th/9608117}}.

\bibitem{Maldacena:1997re}
J.~M. Maldacena, ``{The Large N limit of superconformal field theories and
  supergravity},'' \href{http://dx.doi.org/10.1023/A:1026654312961}{{\em Adv.
  Theor. Math. Phys.} {\bf 2} (1998)  231--252},
  \href{http://arxiv.org/abs/hep-th/9711200}{{\tt arXiv:hep-th/9711200}}.

\bibitem{Gubser:1998bc}
S.~S. Gubser, I.~R. Klebanov, and A.~M. Polyakov, ``{Gauge theory correlators
  from noncritical string theory},''
  \href{http://dx.doi.org/10.1016/S0370-2693(98)00377-3}{{\em Phys. Lett. B}
  {\bf 428} (1998)  105--114}, \href{http://arxiv.org/abs/hep-th/9802109}{{\tt
  arXiv:hep-th/9802109}}.

\bibitem{Witten:1998qj}
E.~Witten, ``{Anti-de Sitter space and holography},''
  \href{http://dx.doi.org/10.4310/ATMP.1998.v2.n2.a2}{{\em Adv. Theor. Math.
  Phys.} {\bf 2} (1998)  253--291},
  \href{http://arxiv.org/abs/hep-th/9802150}{{\tt arXiv:hep-th/9802150}}.

\bibitem{Witten:1995im}
E.~Witten, ``{Bound states of strings and p-branes},''
  \href{http://dx.doi.org/10.1016/0550-3213(95)00610-9}{{\em Nucl. Phys. B}
  {\bf 460} (1996)  335--350}, \href{http://arxiv.org/abs/hep-th/9510135}{{\tt
  arXiv:hep-th/9510135}}.

\bibitem{Aharony:2008ug}
O.~Aharony, O.~Bergman, D.~L. Jafferis, and J.~Maldacena, ``{N=6 superconformal
  Chern-Simons-matter theories, M2-branes and their gravity duals},''
  \href{http://dx.doi.org/10.1088/1126-6708/2008/10/091}{{\em JHEP} {\bf 10}
  (2008)  091}, \href{http://arxiv.org/abs/0806.1218}{{\tt arXiv:0806.1218
  [hep-th]}}.

\bibitem{Duff:2008pa}
M.~J. Duff, ``{Near-horizon brane-scan revived},''
  \href{http://dx.doi.org/10.1016/j.nuclphysb.2008.11.001}{{\em Nucl. Phys. B}
  {\bf 810} (2009)  193--209}, \href{http://arxiv.org/abs/0804.3675}{{\tt
  arXiv:0804.3675 [hep-th]}}.

\bibitem{Gueven:2003uw}
R.~Gueven, J.~T. Liu, C.~N. Pope, and E.~Sezgin, ``{Fine tuning and
  six-dimensional gauged N=(1,0) supergravity vacua},''
  \href{http://dx.doi.org/10.1088/0264-9381/21/4/019}{{\em Class. Quant. Grav.}
  {\bf 21} (2004)  1001--1014}, \href{http://arxiv.org/abs/hep-th/0306201}{{\tt
  arXiv:hep-th/0306201}}.

\bibitem{Edelstein:2010sx}
J.~D. Edelstein, A.~Garbarz, O.~Miskovic, and J.~Zanelli, ``{Stable p-branes in
  Chern-Simons AdS supergravities},''
  \href{http://dx.doi.org/10.1103/PhysRevD.82.044053}{{\em Phys. Rev. D} {\bf
  82} (2010)  044053}, \href{http://arxiv.org/abs/1006.3753}{{\tt
  arXiv:1006.3753 [hep-th]}}.

\bibitem{Apruzzi:2013yva}
F.~Apruzzi, M.~Fazzi, D.~Rosa, and A.~Tomasiello, ``{All AdS$_7$ solutions of
  type II supergravity},''
  \href{http://dx.doi.org/10.1007/JHEP04(2014)064}{{\em JHEP} {\bf 04} (2014)
  064}, \href{http://arxiv.org/abs/1309.2949}{{\tt arXiv:1309.2949 [hep-th]}}.

\bibitem{Gaiotto:2014lca}
D.~Gaiotto and A.~Tomasiello, ``{Holography for (1,0) theories in six
  dimensions},'' \href{http://dx.doi.org/10.1007/JHEP12(2014)003}{{\em JHEP}
  {\bf 12} (2014)  003}, \href{http://arxiv.org/abs/1404.0711}{{\tt
  arXiv:1404.0711 [hep-th]}}.

\bibitem{Cremonesi:2015bld}
S.~Cremonesi and A.~Tomasiello, ``{6d holographic anomaly match as a continuum
  limit},'' \href{http://dx.doi.org/10.1007/JHEP05(2016)031}{{\em JHEP} {\bf
  05} (2016)  031}, \href{http://arxiv.org/abs/1512.02225}{{\tt
  arXiv:1512.02225 [hep-th]}}.

\bibitem{Cordova:2016emh}
C.~Cordova, T.~T. Dumitrescu, and K.~Intriligator, ``{Multiplets of
  Superconformal Symmetry in Diverse Dimensions},''
  \href{http://dx.doi.org/10.1007/JHEP03(2019)163}{{\em JHEP} {\bf 03} (2019)
  163}, \href{http://arxiv.org/abs/1612.00809}{{\tt arXiv:1612.00809
  [hep-th]}}.

\bibitem{Nunez:2018ags}
C.~Nunez, J.~M. Penin, D.~Roychowdhury, and J.~Van~Gorsel, ``{The
  non-Integrability of Strings in Massive Type IIA and their Holographic
  duals},'' \href{http://dx.doi.org/10.1007/JHEP06(2018)078}{{\em JHEP} {\bf
  06} (2018)  078}, \href{http://arxiv.org/abs/1802.04269}{{\tt
  arXiv:1802.04269 [hep-th]}}.

\bibitem{Brandhuber:1999np}
A.~Brandhuber and Y.~Oz, ``{The D-4 - D-8 brane system and five-dimensional
  fixed points},'' \href{http://dx.doi.org/10.1016/S0370-2693(99)00763-7}{{\em
  Phys. Lett. B} {\bf 460} (1999)  307--312},
  \href{http://arxiv.org/abs/hep-th/9905148}{{\tt arXiv:hep-th/9905148}}.

\bibitem{Bergman:2012kr}
O.~Bergman and D.~Rodriguez-Gomez, ``{5d quivers and their AdS(6) duals},''
  \href{http://dx.doi.org/10.1007/JHEP07(2012)171}{{\em JHEP} {\bf 07} (2012)
  171}, \href{http://arxiv.org/abs/1206.3503}{{\tt arXiv:1206.3503 [hep-th]}}.

\bibitem{Lozano:2012au}
Y.~Lozano, E.~\'O~Colg\'ain, D.~Rodriguez-G\'omez, and K.~Sfetsos,
  ``{Supersymmetric $AdS_6$ via T Duality},''
  \href{http://dx.doi.org/10.1103/PhysRevLett.110.231601}{{\em Phys. Rev.
  Lett.} {\bf 110} (2013) no.~23, 231601},
  \href{http://arxiv.org/abs/1212.1043}{{\tt arXiv:1212.1043 [hep-th]}}.

\bibitem{DHoker:2016ujz}
E.~D'Hoker, M.~Gutperle, A.~Karch, and C.~F. Uhlemann, ``{Warped $AdS_6\times
  S^2$ in Type IIB supergravity I: Local solutions},''
  \href{http://dx.doi.org/10.1007/JHEP08(2016)046}{{\em JHEP} {\bf 08} (2016)
  046}, \href{http://arxiv.org/abs/1606.01254}{{\tt arXiv:1606.01254
  [hep-th]}}.

\bibitem{DHoker:2016ysh}
E.~D'Hoker, M.~Gutperle, and C.~F. Uhlemann, ``{Holographic duals for
  five-dimensional superconformal quantum field theories},''
  \href{http://dx.doi.org/10.1103/PhysRevLett.118.101601}{{\em Phys. Rev.
  Lett.} {\bf 118} (2017) no.~10, 101601},
  \href{http://arxiv.org/abs/1611.09411}{{\tt arXiv:1611.09411 [hep-th]}}.

\bibitem{DHoker:2017zwj}
E.~D'Hoker, M.~Gutperle, and C.~F. Uhlemann, ``{Warped $AdS_6\times S^2$ in
  Type IIB supergravity III: Global solutions with seven-branes},''
  \href{http://dx.doi.org/10.1007/JHEP11(2017)200}{{\em JHEP} {\bf 11} (2017)
  200}, \href{http://arxiv.org/abs/1706.00433}{{\tt arXiv:1706.00433
  [hep-th]}}.

\bibitem{Corbino:2017tfl}
D.~Corbino, E.~D'Hoker, and C.~F. Uhlemann, ``{$AdS_{2} \times S^{6}$ versus
  $AdS_{6} \times S^{2}$ in Type IIB supergravity},''
  \href{http://dx.doi.org/10.1007/JHEP03(2018)120}{{\em JHEP} {\bf 03} (2018)
  120}, \href{http://arxiv.org/abs/1712.04463}{{\tt arXiv:1712.04463
  [hep-th]}}.

\bibitem{Lozano:2018pcp}
Y.~Lozano, N.~T. Macpherson, and J.~Montero, ``{AdS$_{6}$ T-duals and type IIB
  AdS$_{6} \times$ S$^{2}$ geometries with 7-branes},''
  \href{http://dx.doi.org/10.1007/JHEP01(2019)116}{{\em JHEP} {\bf 01} (2019)
  116}, \href{http://arxiv.org/abs/1810.08093}{{\tt arXiv:1810.08093
  [hep-th]}}.

\bibitem{Gaiotto:2009gz}
D.~Gaiotto and J.~Maldacena, ``{The Gravity duals of N=2 superconformal field
  theories},'' \href{http://dx.doi.org/10.1007/JHEP10(2012)189}{{\em JHEP} {\bf
  10} (2012)  189}, \href{http://arxiv.org/abs/0904.4466}{{\tt arXiv:0904.4466
  [hep-th]}}.

\bibitem{ReidEdwards:2010qs}
R.~Reid-Edwards and j.~Stefanski, B., ``{On Type IIA geometries dual to N = 2
  SCFTs},'' \href{http://dx.doi.org/10.1016/j.nuclphysb.2011.04.002}{{\em Nucl.
  Phys. B} {\bf 849} (2011)  549--572},
  \href{http://arxiv.org/abs/1011.0216}{{\tt arXiv:1011.0216 [hep-th]}}.

\bibitem{Aharony:2012tz}
O.~Aharony, L.~Berdichevsky, and M.~Berkooz, ``{4d N=2 superconformal linear
  quivers with type IIA duals},''
  \href{http://dx.doi.org/10.1007/JHEP08(2012)131}{{\em JHEP} {\bf 08} (2012)
  131}, \href{http://arxiv.org/abs/1206.5916}{{\tt arXiv:1206.5916 [hep-th]}}.

\bibitem{Nunez:2019gbg}
C.~Nunez, D.~Roychowdhury, S.~Speziali, and S.~Zacarias, ``{Holographic aspects
  of four dimensional $\mathcal{N}=2$ SCFTs and their marginal deformations},''
  \href{http://dx.doi.org/10.1016/j.nuclphysb.2019.114617}{{\em Nucl. Phys. B}
  {\bf 943} (2019)  114617}, \href{http://arxiv.org/abs/1901.02888}{{\tt
  arXiv:1901.02888 [hep-th]}}.

\bibitem{Schwarz:2004yj}
J.~H. Schwarz, ``{Superconformal Chern-Simons theories},''
  \href{http://dx.doi.org/10.1088/1126-6708/2004/11/078}{{\em JHEP} {\bf 11}
  (2004)  078}, \href{http://arxiv.org/abs/hep-th/0411077}{{\tt
  arXiv:hep-th/0411077}}.

\bibitem{DHoker:2007zhm}
E.~D'Hoker, J.~Estes, and M.~Gutperle, ``{Exact half-BPS Type IIB interface
  solutions. I. Local solution and supersymmetric Janus},''
  \href{http://dx.doi.org/10.1088/1126-6708/2007/06/021}{{\em JHEP} {\bf 06}
  (2007)  021}, \href{http://arxiv.org/abs/0705.0022}{{\tt arXiv:0705.0022
  [hep-th]}}.

\bibitem{DHoker:2008lup}
E.~D'Hoker, J.~Estes, M.~Gutperle, and D.~Krym, ``{Exact Half-BPS Flux
  Solutions in M-theory. I: Local Solutions},''
  \href{http://dx.doi.org/10.1088/1126-6708/2008/08/028}{{\em JHEP} {\bf 08}
  (2008)  028}, \href{http://arxiv.org/abs/0806.0605}{{\tt arXiv:0806.0605
  [hep-th]}}.

\bibitem{Chiodaroli:2011nr}
M.~Chiodaroli, E.~D'Hoker, Y.~Guo, and M.~Gutperle, ``{Exact half-BPS
  string-junction solutions in six-dimensional supergravity},''
  \href{http://dx.doi.org/10.1007/JHEP12(2011)086}{{\em JHEP} {\bf 12} (2011)
  086}, \href{http://arxiv.org/abs/1107.1722}{{\tt arXiv:1107.1722 [hep-th]}}.

\bibitem{Assel:2011xz}
B.~Assel, C.~Bachas, J.~Estes, and J.~Gomis, ``{Holographic Duals of D=3 N=4
  Superconformal Field Theories},''
  \href{http://dx.doi.org/10.1007/JHEP08(2011)087}{{\em JHEP} {\bf 08} (2011)
  087}, \href{http://arxiv.org/abs/1106.4253}{{\tt arXiv:1106.4253 [hep-th]}}.

\bibitem{Haupt:2017bnj}
A.~S. Haupt, S.~Lautz, and G.~Papadopoulos, ``{AdS$_{4}$ backgrounds with N
  \ensuremath{>} 16 supersymmetries in 10 and 11 dimensions},''
  \href{http://dx.doi.org/10.1007/JHEP01(2018)087}{{\em JHEP} {\bf 01} (2018)
  087}, \href{http://arxiv.org/abs/1711.08280}{{\tt arXiv:1711.08280
  [hep-th]}}.

\bibitem{Marchesano:2020qvg}
F.~Marchesano, E.~Palti, J.~Quirant, and A.~Tomasiello, ``{On supersymmetric
  AdS$_{4}$ orientifold vacua},''
  \href{http://dx.doi.org/10.1007/JHEP08(2020)087}{{\em JHEP} {\bf 08} (2020)
  087}, \href{http://arxiv.org/abs/2003.13578}{{\tt arXiv:2003.13578
  [hep-th]}}.

\bibitem{Bandres:2008ry}
M.~A. Bandres, A.~E. Lipstein, and J.~H. Schwarz, ``{Studies of the ABJM Theory
  in a Formulation with Manifest SU(4) R-Symmetry},''
  \href{http://dx.doi.org/10.1088/1126-6708/2008/09/027}{{\em JHEP} {\bf 09}
  (2008)  027}, \href{http://arxiv.org/abs/0807.0880}{{\tt arXiv:0807.0880
  [hep-th]}}.

\bibitem{Lozano:2016kum}
Y.~Lozano and C.~N\'u\~nez, ``{Field theory aspects of non-Abelian T-duality
  and $ \mathcal{N} =$ 2 linear quivers},''
  \href{http://dx.doi.org/10.1007/JHEP05(2016)107}{{\em JHEP} {\bf 05} (2016)
  107}, \href{http://arxiv.org/abs/1603.04440}{{\tt arXiv:1603.04440
  [hep-th]}}.

\bibitem{Lozano:2016wrs}
Y.~Lozano, N.~T. Macpherson, J.~Montero, and C.~Nunez, ``{Three-dimensional $
  \mathcal{N}=4 $ linear quivers and non-Abelian T-duals},''
  \href{http://dx.doi.org/10.1007/JHEP11(2016)133}{{\em JHEP} {\bf 11} (2016)
  133}, \href{http://arxiv.org/abs/1609.09061}{{\tt arXiv:1609.09061
  [hep-th]}}.

\bibitem{Couzens:2017way}
C.~Couzens, C.~Lawrie, D.~Martelli, S.~Schafer-Nameki, and J.-M. Wong,
  ``{F-theory and AdS$_{3}$/CFT$_{2}$},''
  \href{http://dx.doi.org/10.1007/JHEP08(2017)043}{{\em JHEP} {\bf 08} (2017)
  043}, \href{http://arxiv.org/abs/1705.04679}{{\tt arXiv:1705.04679
  [hep-th]}}.

\bibitem{Lozano:2017ole}
Y.~Lozano, C.~Nunez, and S.~Zacarias, ``{BMN Vacua, Superstars and Non-Abelian
  T-duality},'' \href{http://dx.doi.org/10.1007/JHEP09(2017)008}{{\em JHEP}
  {\bf 09} (2017)  008}, \href{http://arxiv.org/abs/1703.00417}{{\tt
  arXiv:1703.00417 [hep-th]}}.

\bibitem{Itsios:2017cew}
G.~Itsios, Y.~Lozano, J.~Montero, and C.~Nunez, ``{The AdS$_{5}$ non-Abelian
  T-dual of Klebanov-Witten as a $ \mathcal{N}=1 $ linear quiver from
  M5-branes},'' \href{http://dx.doi.org/10.1007/JHEP09(2017)038}{{\em JHEP}
  {\bf 09} (2017)  038}, \href{http://arxiv.org/abs/1705.09661}{{\tt
  arXiv:1705.09661 [hep-th]}}.

\bibitem{Lozano:2019emq}
Y.~Lozano, N.~T. Macpherson, C.~Nunez, and A.~Ramirez, ``{AdS$_3$ solutions in
  Massive IIA with small $\mathcal{N}=(4,0)$ supersymmetry},''
  \href{http://dx.doi.org/10.1007/JHEP01(2020)129}{{\em JHEP} {\bf 01} (2020)
  129}, \href{http://arxiv.org/abs/1908.09851}{{\tt arXiv:1908.09851
  [hep-th]}}.

\bibitem{Lozano:2019jza}
Y.~Lozano, N.~T. Macpherson, C.~Nunez, and A.~Ramirez, ``{1/4 BPS solutions and
  the AdS$_3$/CFT$_2$ correspondence},''
  \href{http://dx.doi.org/10.1103/PhysRevD.101.026014}{{\em Phys. Rev. D} {\bf
  101} (2020) no.~2, 026014}, \href{http://arxiv.org/abs/1909.09636}{{\tt
  arXiv:1909.09636 [hep-th]}}.

\bibitem{Lozano:2019zvg}
Y.~Lozano, N.~T. Macpherson, C.~Nunez, and A.~Ramirez, ``{Two dimensional
  ${\cal N}=(0,4)$ quivers dual to AdS$_3$ solutions in massive IIA},''
  \href{http://dx.doi.org/10.1007/JHEP01(2020)140}{{\em JHEP} {\bf 01} (2020)
  140}, \href{http://arxiv.org/abs/1909.10510}{{\tt arXiv:1909.10510
  [hep-th]}}.

\bibitem{Lozano:2019ywa}
Y.~Lozano, N.~T. Macpherson, C.~Nunez, and A.~Ramirez, ``{AdS$_3$ solutions in
  massive IIA, defect CFTs and T-duality},''
  \href{http://dx.doi.org/10.1007/JHEP12(2019)013}{{\em JHEP} {\bf 12} (2019)
  013}, \href{http://arxiv.org/abs/1909.11669}{{\tt arXiv:1909.11669
  [hep-th]}}.

\bibitem{Lozano:2020bxo}
Y.~Lozano, C.~Nunez, A.~Ramirez, and S.~Speziali, ``{$M$-strings and AdS$_3$
  solutions to M-theory with small $\mathcal{N}=(0,4)$ supersymmetry},''
  \href{http://dx.doi.org/10.1007/JHEP08(2020)118}{{\em JHEP} {\bf 08} (2020)
  118}, \href{http://arxiv.org/abs/2005.06561}{{\tt arXiv:2005.06561
  [hep-th]}}.

\bibitem{Faedo:2020nol}
F.~Faedo, Y.~Lozano, and N.~Petri, ``{Searching for surface defect CFTs within
  AdS$_3$},'' \href{http://dx.doi.org/10.1007/JHEP11(2020)052}{{\em JHEP} {\bf
  11} (2020)  052}, \href{http://arxiv.org/abs/2007.16167}{{\tt
  arXiv:2007.16167 [hep-th]}}.

\bibitem{Faedo:2020lyw}
F.~Faedo, Y.~Lozano, and N.~Petri, ``{New $\mathcal{N}=(0,4)$ AdS$_3$
  near-horizons in Type IIB},'' \href{http://arxiv.org/abs/2012.07148}{{\tt
  arXiv:2012.07148 [hep-th]}}.

\bibitem{Dibitetto:2020bsh}
G.~Dibitetto and N.~Petri, ``{AdS$_3$ from M-branes at conical
  singularities},'' \href{http://arxiv.org/abs/2010.12323}{{\tt
  arXiv:2010.12323 [hep-th]}}.

\bibitem{Macpherson:2021lbr}
N.~T. Macpherson and A.~Tomasiello, ``{${\cal N}=(1,1)$ supersymmetric AdS$_3$
  in 10 dimensions},'' \href{http://arxiv.org/abs/2110.01627}{{\tt
  arXiv:2110.01627 [hep-th]}}.

\bibitem{Dibitetto:2018gtk}
G.~Dibitetto and N.~Petri, ``{AdS$_{2}$ solutions and their massive IIA
  origin},'' \href{http://dx.doi.org/10.1007/JHEP05(2019)107}{{\em JHEP} {\bf
  05} (2019)  107}, \href{http://arxiv.org/abs/1811.11572}{{\tt
  arXiv:1811.11572 [hep-th]}}.

\bibitem{Gauntlett:2006ns}
J.~P. Gauntlett, N.~Kim, and D.~Waldram, ``{Supersymmetric AdS(3), AdS(2) and
  Bubble Solutions},''
  \href{http://dx.doi.org/10.1088/1126-6708/2007/04/005}{{\em JHEP} {\bf 04}
  (2007)  005}, \href{http://arxiv.org/abs/hep-th/0612253}{{\tt
  arXiv:hep-th/0612253}}.

\bibitem{Kim:2013xza}
N.~Kim, ``{Comments on $AdS_2$ solutions from M2-branes on complex curves and
  the backreacted K$\ddot{\text{a}}$hler geometry},''
  \href{http://dx.doi.org/10.1140/epjc/s10052-014-2778-6}{{\em Eur. Phys. J. C}
  {\bf 74} (2014) no.~2, 2778}, \href{http://arxiv.org/abs/1311.7372}{{\tt
  arXiv:1311.7372 [hep-th]}}.

\bibitem{Chiodaroli:2009yw}
M.~Chiodaroli, M.~Gutperle, and D.~Krym, ``{Half-BPS Solutions locally
  asymptotic to AdS(3) x S**3 and interface conformal field theories},''
  \href{http://dx.doi.org/10.1007/JHEP02(2010)066}{{\em JHEP} {\bf 02} (2010)
  066}, \href{http://arxiv.org/abs/0910.0466}{{\tt arXiv:0910.0466 [hep-th]}}.

\bibitem{Chiodaroli:2009xh}
M.~Chiodaroli, E.~D'Hoker, and M.~Gutperle, ``{Open Worldsheets for Holographic
  Interfaces},'' \href{http://dx.doi.org/10.1007/JHEP03(2010)060}{{\em JHEP}
  {\bf 03} (2010)  060}, \href{http://arxiv.org/abs/0912.4679}{{\tt
  arXiv:0912.4679 [hep-th]}}.

\bibitem{Corbino:2018fwb}
D.~Corbino, E.~D'Hoker, J.~Kaidi, and C.~F. Uhlemann, ``{Global half-BPS
  $AdS_{2} \times S^6$ solutions in Type IIB},''
  \href{http://dx.doi.org/10.1007/JHEP03(2019)039}{{\em JHEP} {\bf 03} (2019)
  039}, \href{http://arxiv.org/abs/1812.10206}{{\tt arXiv:1812.10206
  [hep-th]}}.

\bibitem{Corbino:2020lzq}
D.~Corbino, ``{Warped $AdS_{2}$ and $SU(1,1|4)$ symmetry in Type IIB},''
  \href{http://arxiv.org/abs/2004.12613}{{\tt arXiv:2004.12613 [hep-th]}}.

\bibitem{Dibitetto:2019nyz}
G.~Dibitetto, Y.~Lozano, N.~Petri, and A.~Ramirez, ``{Holographic description
  of M-branes via AdS$_{2}$},''
  \href{http://dx.doi.org/10.1007/JHEP04(2020)037}{{\em JHEP} {\bf 04} (2020)
  037}, \href{http://arxiv.org/abs/1912.09932}{{\tt arXiv:1912.09932
  [hep-th]}}.

\bibitem{Lozano:2020txg}
Y.~Lozano, C.~Nunez, A.~Ramirez, and S.~Speziali, ``{New AdS$_2$ backgrounds
  and ${\cal N}=4$ Conformal Quantum Mechanics},''
  \href{http://arxiv.org/abs/2011.00005}{{\tt arXiv:2011.00005 [hep-th]}}.

\bibitem{Lozano:2020sae}
Y.~Lozano, C.~Nunez, A.~Ramirez, and S.~Speziali, ``{AdS$_2$ duals to ADHM
  quivers with Wilson lines},'' \href{http://arxiv.org/abs/2011.13932}{{\tt
  arXiv:2011.13932 [hep-th]}}.

\bibitem{DHoker:2007mci}
E.~D'Hoker, J.~Estes, and M.~Gutperle, ``{Gravity duals of half-BPS Wilson
  loops},'' \href{http://dx.doi.org/10.1088/1126-6708/2007/06/063}{{\em JHEP}
  {\bf 06} (2007)  063}, \href{http://arxiv.org/abs/0705.1004}{{\tt
  arXiv:0705.1004 [hep-th]}}.

\bibitem{Bachas:2013vza}
C.~Bachas, E.~D'Hoker, J.~Estes, and D.~Krym, ``{M-theory Solutions Invariant
  under $D(2,1;\gamma) \oplus D(2,1;\gamma)$},''
  \href{http://dx.doi.org/10.1002/prop.201300039}{{\em Fortsch. Phys.} {\bf 62}
  (2014)  207--254}, \href{http://arxiv.org/abs/1312.5477}{{\tt arXiv:1312.5477
  [hep-th]}}.

\end{thebibliography}\endgroup

\documentclass[12pt,a4paper
]{article}
\usepackage{amsmath, amsthm,mathtools, amssymb,upgreek,slashed,hyperref}
\usepackage[utf8]{inputenc}
\usepackage{ifpdf}
\ifpdf
  \usepackage[pdftex]{graphicx}
  \usepackage{xcolor}
  \usepackage{epstopdf}
  \usepackage{bigints}
\else
  \usepackage[dvips]{graphicx}
\fi
\usepackage[paper=letterpaper,margin=0.9in]{geometry}
\usepackage{tikz}
\def\Bbb{\mathbb}

\def\SU{{\mathrm{\SU}}}

\def\OSp{{\textrm{OSp}}}
\def\SU{{\textrm{SU}}}
\def\SO{{\textrm{SO}}}
\def\USp{{\textrm{USp}}}
\def\Sp{{\textrm{Sp}}}
\def\Spin{{\textrm{Spin}}}
\def\Spin{{\textrm{Spin}}}
\def\Or{{\textrm{O}}}
\def\Un{{\textrm{U}}}

\def\U{{\mathrm U}}

\def\2{{\bf 2}}
\def\3{{\bf 3}}

\def\4{{\bf 4}}

\def\be{\begin{equation}}
\def\ee{\end{equation}}
 \def\\Sp{{\mathrm{\Sp}}}
 \def\\Spin{{\mathrm{\Spin}}}
 \def\\SU{{\mathrm{\SU}}}
 \def\\SO{{\mathrm{\SO}}}

\def\\Sp{\mathrm{\Sp}}

\def\R{{\Bbb{R}}}

\font\teneurm=eurm10 \font\seveneurm=eurm7 \font\fiveeurm=eurm5
\newfam\eurmfam
\textfont\eurmfam=\teneurm \scriptfont\eurmfam=\seveneurm
\scriptscriptfont\eurmfam=\fiveeurm

\font\teneusm=eusm10 \font\seveneusm=eusm7 \font\fiveeusm=eusm5
\newfam\eusmfam
\textfont\eusmfam=\teneusm \scriptfont\eusmfam=\seveneusm
\scriptscriptfont\eusmfam=\fiveeusm

\font\tencmmib=cmmib10 \skewchar\tencmmib='177
\font\sevencmmib=cmmib7 \skewchar\sevencmmib='177
\font\fivecmmib=cmmib5 \skewchar\fivecmmib='177
\newfam\cmmibfam
\textfont\cmmibfam=\tencmmib \scriptfont\cmmibfam=\sevencmmib
\scriptscriptfont\cmmibfam=\fivecmmib

\numberwithin{equation}{section}

\def\C{{\Bbb C}}

\def\\OSp{{\mathrm{\OSp}}}

\vspace{10pt}

\begin{document}
\begin{titlepage}
\begin{flushright}
\hfill{Imperial/TP/2021/mjd/3}\\

%\hfill{hep-th/yymmnnn}\\
\end{flushright}
\begin{center}

\vskip .5in %.3in 
\noindent

{\Large \bf{ The conformal brane-scan: an update }}
\bigskip\medskip

M. J. Duff\footnote{m.duff@imperial.ac.uk},\\
\bigskip\medskip

{\small

 Department of Physics, Imperial College London, Prince Consort Road, London SW7 2BZ
\vskip 3mm
\begin{center}
\&
\end{center}
Institute for Quantum Science and Engineering and Hagler Institute for Advanced Study, Texas A\&M University, College Station, TX, 77840, USA
}

\vskip .5cm %.3cm
\vskip .9cm %.6cm
     	{\bf Abstract }

\vskip .1in
\end{center}

\noindent

Generalizing the {\it The Membrane at the End of the Universe}, a 1987 paper {\it Supersingletons} by Blencowe and the author conjectured the existence of BPS $p$-brane configurations ($p=2, 3, 4 ,5$) and corresponding CFTs on the boundary of anti-de Sitter space with symmetries appearing in Nahm's classification of superconformal algebras: ${ \OSp(N|4)}~N=8, 4, 2, 1$; ${ \SU(2,2|N)}~N= 4, 2, 1$; 
${F^2(4)}$; ${\OSp(8^*|N)}, ~N=4, 2$. This correctly predicted the $D3$-brane with ${ \SU(2,2|4)}$ on $AdS_5 \times S^5$ and the $M5$-brane with ${ \OSp(8^*{}|4)}$ on $AdS_7 \times S^4$,  in addition to the known $M2$-brane with ${ \OSp(8|4)}$ on $AdS_4 \times S^7$.  However, finding non-singular AdS solutions matching the other symmetries was less straightforward.  Here we perform a literature search and confirm that all of the empty slots have now been filled, thanks to a number of extra ingredients including warped products and massive Type IIA. Orbifolds, orientifolds and S-folds also play a part providing examples not predicted: ${ \SU(2,2|3)}$, ${ \OSp(3|4)}$, ${ \OSp(5|4)}$ and ${ \OSp(6|4)}$ but not  ${ \OSp(7|4)}$.  We also examine the status of $p=(0,1)$ configurations.

\noindent
\vskip .5cm
\vskip .5cm

\vfill
\eject

\end{titlepage}

\setcounter{footnote}{0}
\setcounter{tocdepth}{2}

{\parskip = .2\baselineskip \tableofcontents}
\newpage
%\section{Introduction}

%\section{\bf Two brane-scans}
{\it Our mistake is not that we take our theories too seriously, but that we do not take them seriously enough.}

  Steven Weinberg

\section{Supersingletons}

\indent

The {\it Membrane at the End of the Universe} \cite{Duff:1987qa,Bergshoeff:1987dh,Blencowe:1987bn,Duff:1988st,Nicolai:1988ek,Bergshoeff:1988uc,Duff:1989ez,
Bergshoeff:1989av,Duff:1989aj,Claus:1998mw} was the name given to a supermembrane  \cite{Bergshoeff:1987cm} (later called the M2-brane) on  the $S^1 \times S^2$ boundary
 of $AdS_4 \times S^7$ described by a SCFT with symmetry 
\be
\OSp(8|4) \supset \SO(3,2) \times \SO(8) 
\ee
namely the $N=8$ singleton supermultiplet with 8 scalar and 8 spinors and $\SO(8)$ $R$ symmetry.  We recall that representations of $\SO(3,2)$ are denoted $D(E_0,s)$ where $E_0$ is the lowest energy eigenvalue which occurs and $s$ is the total angular momentum quantum number of the lowest energy state, analogous to the mass and spin of the Poincare group. However, Dirac's singletons $D(1/2,0)$ and $D(1,1/2)$ have no four-dimensional Poincare analogue \cite{Dirac:1963ta} and are best interpreted a residing on the three-dimensional boundary \cite{Fronsdal:1981gq,Nicolai:1984gb,Bergshoeff:1987dh}.
\begin{table}[h]
\small{
\begin{center}
\begin{tabular}{llllllllllllllllllllllllll}
d&G&H&&&Susy\\\\
6&$\OSp(8^{*}|N)$&$\SO^{*}(8) \times \USp(N)$&$N~ even$&&$8N$\\
5&$F^2{}(4)$&$ \SO(5,2) \times \SU(2)$&&&16\\
4&$\SU(2,2|N)$&$\SU(2,2) \times \Un(N)$&$N\neq 4$&&$8N$\\
~&$\SU(2,2|4)$&$\SU(2,2)\times \SU(4)$&&&32\\
3&$\OSp(N|4)$&$ \SO(N) \times \Sp(4, \R)$&&&$4N$\\
2&$G_{+} \times G_{-}$\\
1&$G_{\pm}=$\\
&$\OSp(N|2)$&$\Or(N) \times \SU(1,1)$&&&$2N$\\
& $\SU(N|1,1)$ & $\Un(N) \times \SU(1,1)$&$N \neq 2 $&&$4N$\\
&$\SU(2|1,1) $&$\SU(2) \times \SU(1,1)$&&&8 \\
&$\OSp(4^{*}|2N)$ &$  \SU(2) \times \U\Sp(2N) \times \SU(1,1)$&&&$8N$\\
&$G(3) $&$G_{2} \times \SU(1,1)$&&&14\\
&$F(4) $&$\Spin(7) \times \SU(1,1)$&&&16\\
&$D^{1}(2,1,\alpha) $&$\SU(2) \times \SU(2) \times \SU(1,1)$&&&8 \\
\end{tabular}
\end{center}}
\caption{Following \cite{Nahm:1977tg,VanProeyen:1999ni} we list the AdS supergroups in $d\le 6$ and their bosonic subgroups in the notation of \cite{Gunaydin:1998km}.}
\label{Nahm}
\end{table}
% with bosonic subgroups $\SO(p+1,2) \times \SO(D-p-1)$ describing $p$-branes on the boundary of  $AdS_{p+2} \times S^{D-p-2}$, as shown in Table \ref{horscan}. 

\begin{table}[h]
\begin{center}
\begin{tabular}{ccccccccccccccc}
&{\bf Supergroup}&{\bf Supermultiplet}&$B^-$&$V$&$\chi$&$\phi$&&$D$&\\\\
$AdS_3$&$\OSp(n|2) \times \OSp(8-n|2)$&$(n_+,n_-)=(n,8-n),d=2$~singleton&0&0&8&8&&10\\
&$\OSp(n|2) \times \OSp(4-n|2)$&$(n_+,n_-)=(n,4-n),d=2$~singleton&0&0&4&4&&6\\
&$\OSp(n|2) \times \OSp(2-n|2)$&$(n_+,n_-)=(n,2-n),d=2$~singleton&0&0&2&2&&4\\
&$\OSp(n|2) \times \OSp(1-n|2)$&$(n_+,n_-)=(n,1-n),d=2$~singleton&0&0&1&1&&3\\\\
$AdS_4$&$\OSp(8|4)$&$n=8,d=3$~singleton&0&0&8&8&&11&\\
&$\OSp(4|4)$&$n=4,d=3$~singleton&0&0&4&4&&7&\\
&$\OSp(2|4)$&$n=2,d=3$~singleton&0&0&2&2&&5&\\
&$\OSp(1|4)$&$n=1,d=3$~singleton&0&0&1&1&&4&\\\\
$AdS_5$&$\SU(2,2|2)$&$n=2,d=4$~doubleton&0&0&2&4&&8&\\
&$\SU(2,2|1)$&$n=1,d=4$~doubleton&0&0&1&2&&6&\\\\
&$\SU(2,2|4)$&$n=4,d=4$~doubleton&0&1&4&6&&10&\\
&$\SU(2,2|2)$&$n=2,d=4$~doubleton&0&1&2&2&&6&\\
&$\SU(2,2|1)$&$n=1,d=4$~doubleton&0&1&1&0&&4&\\\\
$AdS_6$&$F^2(4)$&$n=2,d=5$~doubleton&0&0&2&4&&9&\\\\
$AdS_7$&$\OSp(8^*|2)$&$(n_{+},n_{-})=(1,0),d=6$~tripleton&0&0&1&4&&10\\\\
&$\OSp(8^*|4)$&$(n_{+},n_{-})=(2,0),d=6$~tripleton&1&0&2&5&&11&\\
&$\OSp(8^*|2)$&$(n_{+},n_{-})=(1,0),d=6$~tripleton&1&0&1&1&&7&\\
\end{tabular}
\\
\end{center}
\medskip
\caption{Superconformal groups and their singleton, doubleton 
and tripleton representations. $B^-$, $V$, $\chi$, $\phi$ denote the number of chiral 2-forms, vector, spinors and scalars in each multiplet.  The spacetime dimension $D$ equals the worldvolume dimension $d$ plus the number of scalars.}
\label{fields}
\end{table}

Accordingly, in 1987
Blencowe and the author \cite{Blencowe:1987bn} conjectured the existence of  other BPS $p$-brane configurations with $p=(2,3,4,5)$ on the $S^1 \times S^p$ boundary of $AdS_{(p+2) }$ and corresponding CFTs with other symmetries appearing in Nahm's classification of superconformal algebras \cite{Nahm:1977tg}, listed in Table \ref{Nahm}. 

In each case the boundary CFT is described by the corresponding singleton (scalar), doubleton (scalar or vector) or tripleton (scalar or tensor) supermultiplet\footnote{Our nomenclature, based on the rank of $AdS_{p+2}$, is singleton $p=2$, doubleton $p=(2,3)$, tripleton $p=5$ and differs from that of G\"{u}naydin and Minic \cite{Gunaydin:1998km}.} as shown in Table \ref{fields}.  The number of dimensions transverse to the brane, $D-d$, equals the number of scalars in the supermultiplets.  None of these BPS brane CFTs is self-interacting. (For non-BPS see \cite{Seiberg:1999xz,Batrachenko:2002pu}).

\begin{table}
\small{
\begin{center}
\begin{tabular}{ccccccccccccccccc}
D$\uparrow$\\\\
%\hline\\
&&~{\bf SCALAR}\\\
11&.&&&${\bf OSp(8|4)}$&&&\\\
10&.&&$\OSp(n|2) \times \OSp(8-n|2)$&&&&$\OSp(8^*|2)$\\\
9&.&&&&&$F^2(4)$&\\\
8&.&&&&$\SU(2,2|2)$&&\\\
7&.&&&$\OSp(4|4)$&&&\\\
6&.&&$\OSp(n|2) \times \OSp(4-n|2)$&&$\SU(2,2|1)$\\\
5&.&&&$\OSp(2|4)$&&&\\\
4&.&&$\OSp(n|2) \times \OSp(2-n|2)$&$\OSp(1|4)$&&&\\\
3&.&&$\OSp(n|2) \times \OSp(1-n|2)$&~&&&\\\
2&.&&&&&&&&\\\
1&.&&\\\
0&.&.&.&.&.&.&\\\
&&~{\bf VECTOR}\\\
11&.&\\\
10&.&&&&${\bf SU(2,2|4)}$&&\\\
9&.&&&&&&\\\
8&.&&&&&&\\\
7&.&&&&&\\\
6&.&&&&$\SU(2,2|2)$&&\\\
5&.&\\
4&.&&&&$\SU(2,2|1)$&&\\\
3&.\\\
2&.&&&\\\
1&.&&\\
0&.&.&.&.&.&.&.\\\
&&~{\bf TENSOR}\\\
11&&&&&&&${\bf OSp(8^*|4)}$\\\
10&.&&&&&\\\
9&.&&&&&&\\\
8&.&&&&&&&\\\
7&.&&&&&&$\OSp(8^*|2)$\\\
6&.&&&&\\\
5&.&&&\\\
4&.&&&&&\\\
3&.&&&&&\\\
2&.&&&&&&&\\\
1&.&&&&&&\\\
0&.&.&.&.&.&.&.\\\
&0&1&2&3&4&5&6&d$\rightarrow$
\end{tabular}
\end{center}
\caption{The brane-scans of superconformal groups:
scalar supermultiplets: singletons ($p=1,2$), doubletons ($p=3,4)$ and tripletons ($p=5$);
vector supermultiplets: doubletons ($p=3$); 
tensor supermultiplets: tripletons ($p=5$). 
The M2-, D3- and M5-branes are in boldface.}
\label{tensor}}
\end{table}

A plot of spacetime dimension $D$ vs worldvolume dimension $d=p+1$, known as the {\it brane-scan}, is shown in Table \ref{tensor}.   This correctly predicted the $D3$-brane \cite{Horowitz:1991cd,Duff:1991pea,Duff:1992hu,Duff:1994an,Polchinski:1995mt,Polchinski:1996na} with ${\bf SU(2,2|4)}$ on $AdS_5 \times S^5$ and the $M5$-brane \cite{Gueven:1992hh,Duff:1992hu,Duff:1994an}  with ${\bf OSp(8^*|4)}$ on $AdS_7 \times S^4$,  in addition to the known $M2$-brane \cite{Bergshoeff:1987cm,Duff:1994an} with ${\bf OSp(8|4)}$ on $AdS_4 \times S^7$. 
The purpose of the present paper is to report  that all of the other slots have now been filled, thanks to a number of extra ingredients: warped products, massive Type IIA and Chern-Simons theories. Orbifolds, orientifolds and S-folds also play a part providing examples not predicted: ${ \SU(2,2|3)}$, ${ \OSp(3|4)}$, ${ \OSp(5|4)}$ and ${ \OSp(6|4)}$ but not  ${ \OSp(7|4)}$. We also examine the status of $p=(0,1)$ configurations. 
%\FloatBarrier

\section{The conformal brane-scan}

Comments:
\begin{itemize}
\item

The list in Table \ref{Nahm} is complete if one assumes that the Killing superalgebras of $AdS$ backgrounds are simple. However a more detailed investigation reveals that there may be some additional central generators in the Killing superlgebra for $AdS_3 $ and $AdS_5$ backgrounds \cite{Beck:2017wpm,Haupt:2018gap}

\item
The supersingleton lagrangian and transformation rules were also spelled out explicitly in \cite{Blencowe:1987bn}. This {\it conformal}  or (in later terminology) {\it near-horizon} brane-scan differs from the scan of Green-Schwarz type kappa-symmetric branes \cite{Achucarro:1987nc} which are not in general conformal and which, in any case, include only scalar supermultiplets. 
Further developments and elaborations on the brane-scan are summarized in \href{https://ncatlab.org/nlab/show/brane+scan}{Schreiber's n-lab} 
 and references  therein.
\item
In early 1988, Nicolai, Sezgin and Tanii \cite{Nicolai:1988ek} independently put forward the same generalization of the {\it Membrane at the End of the Universe} idea, spelling out the doubleton and tripleton lagrangian and transformation rules, in addition to the singleton.
However, by insisting on only scalar supermultiplets as in \cite{Achucarro:1987nc}  their list excluded the vector or tensor brane-scans of Table \ref{tensor}. In this case, as they 
point out, the spheres are just the parallelizable ones $S^1$, $S^3$ and $S^7$.

\item
 The two factors appearing in the $p=1$ case, $G_+ \times G_-$, are simply a reflection of the ability of strings to have left and right movers on the worldsheet \cite{Gunaydin:1986cs}.  
In this case, there are many candidate supergroups as shown in Table \ref{Nahm}, so for $p=0,1$ we did not attempt a complete list of which of these would eventually be realized. In  \cite{Blencowe:1987bn}, we focused on Type IIA, Type IIB and heterotic strings
with ${  \OSp(n|2)}_c \times { \OSp(8-n|2)}_s$, $ { \OSp(n|2)}_c \times { \OSp(8-n|2)}_c$ and $ { \OSp(n|2)}_c\times { \Sp(2, \R)}$, respectively, since the singleton CFTs (but not the supergravity $AdS_3$ solutions) had already been identified \cite{Gunaydin:1986cs}.  For concreteness the Type IIA case appears on the scan of Table \ref{tensor}. 

\item
Even for $p \geq 2$ not all  of the conformal algebras listed in Table \ref{Nahm} appear in the scan. For example, since none of our CFTs is self-interacting, we restricted  \cite{Blencowe:1987bn} ${ \SU(2,2|N)} $ to $N=1, 2, 4$ since perturbatively $N=3$ implies $N=4$. But we now know there are nonperturbative interacting CFTs with just $N=3$ \cite{Ferrara:1998zt,Aharony:2015oyb,Garcia-Etxebarria:2015wns,Aharony:2016kai,Borsten:2018jjm}.
We also focussed on $ N=1, 2, 4, 8$ in ${\OSp(N|4)}$ since they corresponded to the division algebra $\R, \C, \mathbb{H}, \mathbb{O}$ interpretation  of the four diagonal lines in the scalar branescan of Table \ref{tensor}. The $N=3,5,6,7$ cases are discussed in Section \ref{missing}.
\end{itemize}

\section{Significance of the brane-scan}

The significance of the $M2$, $D3$ and $M5$ and indeed the other  configurations on the brane-scan became clearer thanks to four major developments:

\begin{itemize}
 \item Branes as solitons 
 
The realization that string theory admits p-branes as solitons \cite{Townsend:1987yy,Strominger:1990et,Duff:1991sz,Duff:1991pe,Duff:1991pea,Horowitz:1991cd,Callan:1991at,Gibbons:1993xt,Duff:1994an}

 \item M-theory
 
  The realization that the Type IIA superstring in $D=10$ could be interpreted \cite{Duff:1987bx} as a wrapped supermembrane in $D=11$ \cite{Bergshoeff:1987cm}.  
The membrane is a 1/2 BPS solution of $D=11$ supergravity \cite{Duff:1990xz}, whose spacetime approaches Minkowski space far away from the brane but  $AdS_4 \times S^7$ close to the brane, jumping to the full ${\OSp(8|4)}$ in the limit \cite{Duff:1994fg}.  Regarded as an extremal black-brane, this limit was also called the near-horizon limit. Moreover multi-brane solutions could be obtained by stacking $N$ branes on top of one another \cite{Duff:1990xz}, yielding  quantized 4-form flux.  So $AdS_4 \times S^7$ could equally well be regarded as the large $N$ limit. A similar story applied to its magnetic dual fivebrane \cite{Gueven:1992hh} as a solution of $D=11$ supergravity. Moreover, the five string theories were merely different corners of an overarching M-theory \cite{Hull:1994ys,Witten:1995ex,Duff:1996aw} with $D=11$ supergravity as its low-energy limit. The membrane and fivebrane  were accordingly renamed M2 and M5.

\item D-branes

The realization that p-branes carrying RR charge, with a closed-string interpretation as solitons, admitted an alternative open string interpretation as Dirichlet-branes, surfaces of dimension $p$ on which open strings can end \cite{Polchinski:1996na}.  In particular the self-dual 3-brane, a solution of Type IIB supergravity with $AdS_5 \times S^5$ and ${ \SU(2,2|4)}$ in the large N limit, was reinterpreted as a D3-brane and renamed accordingly. 

\item AdS/CFT

The AdS/CFT  conjecture \cite{Maldacena:1997re,Gubser:1998bc,Witten:1998qj} proposes that large N limits of certain conformal field theories in d dimensions can be described in terms of supergravity (and string theory) on the product of d+1-dimensional AdS space with a compact manifold. Another vital ingredient, missing in the early days, was the non-abelian nature of the symmetries that appear when we stack N branes on top of one another \cite{Witten:1995im}. Examples include $N=4$ Yang-Mills in $D=4$ from $AdS_5 \times S^5$ and ABJM theory \cite{Aharony:2008ug} from $AdS_4 \times S^7/Z_n$.

\end{itemize}
    
\section{The missing ingredients $p\geq2$}
\label{missing}
Notwithstanding the success with $M2$, $D3$ and $M5$, for quite some time the status of the other slots on the brane-scans remained obscure \footnote{In  \cite{Duff:2008pa} we entertained the idea that they might arise from classical branes whose symmetry is enhanced when $\alpha'$ corrections are taken into account, but this did not pan out.}.
Here we perform a literature search and confirm that  all of the empty slots have now been filled, largely thanks to warped products, massive Type IIA, and Chern Simons theories as shown below

\begin{itemize}
\item{d=6}
$\OSp(8^*|N)$ $N=4,2$; \cite{Gueven:2003uw,Edelstein:2010sx,Apruzzi:2013yva, Gaiotto:2014lca, Cremonesi:2015bld,Cordova:2016emh,Nunez:2018ags} 

\item{d=5} 
$F^2(4)$ \cite{Brandhuber:1999np,Bergman:2012kr,Lozano:2012au,DHoker:2016ujz,DHoker:2016ysh,Cordova:2016emh,DHoker:2017zwj,
Corbino:2017tfl,Lozano:2018pcp}

\item{d=4}
$ \SU(2,2|N)$ $N= 4, 3, 2, 1$; 
\cite{Horowitz:1991cd,Gaiotto:2009gz,ReidEdwards:2010qs, Aharony:2012tz,Cordova:2016emh,Nunez:2019gbg,Ferrara:1998zt,Aharony:2015oyb,Garcia-Etxebarria:2015wns,Aharony:2016kai,Borsten:2018jjm}.

\item{d=3}
$\OSp(N|4)$  $N=8,6,5,4,3,2,1$ \cite{Duff:1990xz,Schwarz:2004yj,DHoker:2007zhm, DHoker:2008lup,Aharony:2008ug,Chiodaroli:2011nr, Assel:2011xz,Cordova:2016emh,Haupt:2017bnj,Marchesano:2020qvg,Bandres:2008ry}.

\end{itemize}
Comments
\begin{itemize}
\item
We have included $N=3$ in the $d=4$ case and $N=3,5,6$ in the $d=3$ case, which, as already mentioned ,were not predicted in  \cite{Blencowe:1987bn}. $N=6$ appears in ABJM \cite{Aharony:2008ug}. and its  ${\OSp(6|4)}$  symmetry in \cite{ Bandres:2008ry}.  A useful reference on the absence of $N=7$ is \cite{Cordova:2016emh}.

\item
There are no $AdS_7$ solutions in Types IIA and IIB. In M all are locally isometric to $AdS_7 \times S^4$.
\item
There are no maximally supersymmetric $AdS_6$ backgrounds in M, IIA or IIB. There are no half BPS (16 supersymmetries) $AdS_6$ backgrounds in M and IIA with compact internal space.
\item
There are no such $AdS_5$ solutions that preserve $>16$ supersymmetries in IIA and D=11
In IIB, all supersymmetric solutions are locally isometric to $AdS_5\times S^5$. This means
that all backgrounds preserving 24 supersymmetries in IIB are locally   $AdS_5\times S^5.$
\item

There are  no $ >16$ $AdS_4$ supersymmetric solutions in IIA and IIB.
In D=11 all  $>16$ supersymmetric solutions are locally isometric to $AdS_4\times S^7$.
This means that all solutions with 20, 24, 28 are locally  $AdS_4\times S^7.$
\end{itemize}

\section{$p=0,1$}
\begin{itemize}

\item{d=2}
\cite{Edelstein:2010sx,Lozano:2016kum,Lozano:2016wrs,Couzens:2017way,Lozano:2017ole,Itsios:2017cew, 
Haupt:2017bnj,Lozano:2019emq,Lozano:2019jza,Lozano:2019zvg,Lozano:2019ywa,Lozano:2020bxo,Faedo:2020nol,Faedo:2020lyw,Dibitetto:2020bsh,Macpherson:2021lbr}

\item{d=1} 
 \cite{Dibitetto:2018gtk, Gauntlett:2006ns, Kim:2013xza, Chiodaroli:2009yw, Chiodaroli:2009xh,Corbino:2018fwb,Corbino:2020lzq,Dibitetto:2019nyz,Lozano:2020txg,Lozano:2020sae,DHoker:2007mci,Bachas:2013vza}

\end{itemize}

Comment
\begin{itemize}
\item
Not all of the algebras in Nahm's list correspond to known solutions and indeed there may be some for which no solutions exist. 
A thorough and up-to-date summary maybe found in \cite{Macpherson:2021lbr}.

\end{itemize}

\section{Conclusion} 
Thus not only the M2, D3 and M5 but all of the $p$-brane configurations on the $S^1 \times S^p$ boundary of $AdS_{(p+1) }$ with $p=(5,4,3,2,1)$ mentioned explicitly in the 1987 paper as shown in Table  3 have now been discovered: ${ \OSp(N|4)}~N=8, 4, 2, 1$; ${\bf \SU(2,2|N)}~N= 4, 2, 1$; 
${ F^2(4)}$; ${ \OSp(8^*|N)}, ~N=4, 2$,  as have most of the $(p=0,1)$ in Nahm's list not mentioned explicitly. Orbifolds, orientifolds and S-folds also play a part providing examples not predicted: ${ \SU(2,2|3)}$, ${ \OSp(3|4)}$, ${ \OSp(5|4)}$ and ${\OSp(6|4)}$ but not  ${ \OSp(7|4)}$.  
To be fair, if our colleagues did not take our vector and tensor brane-scans seriously in 1987, it may be because, in the Weinberg sense, we did not take them seriously enough ourselves. 
\section*{Acknowledgements}
 
Correspondence with Fernando Alday, Connor Behan, Leron Borsten, Eric D'Hoker, George Papadopoulos, Yolanda Lozano and  Alessandro Tomasiello is greatly appreciated. I am grateful to Marlan Scully for his hospitality in the Institute for Quantum Science and Engineering, Texas A\&M University, and to the Hagler Institute for Advanced Study at Texas A\&M for a Faculty Fellowship. 
This work was supported in part by the STFC under rolling grant ST/P000762/1. 
\bibliography{Bibliography}
%\bibliography{references}
%\bibliographystyle{bibliography}   
\end{document}